Iodine molecule dissociates into atoms after absorbing light of 4500 Angstrom.if one quantum of radiation is absorbed by each molecule,calculate the kinetic energy of iodine atoms.(Bond energy of I2 =240 kJ mol-1)

Grade:12th Pass

1 Answers

Rahul Kumar
12 years ago

Enegy of 1 quanta of radiations = hc/wavelength

                                 = 6.6*10^-34 * 3*10^8/4500*10*-10

                                 = 4.4*10^-19

Bond energy per molecule = 240kj/mol

                                     = 240 * 1000/6*10^23

                                     = 4*10^-19

Hence kinetic energy of iodine atoms = 4.4*10^-19 - 4*10^-19